OpenStudy (anonymous):

Find the inverse of the function. f(x) = the cube root of (x/7) -9

OpenStudy (anonymous):

\[\sqrt[3]{(x/7)-9}\] Is that what yu mean?

OpenStudy (anonymous):

Yes exactly :)

OpenStudy (anonymous):

Write f(x) as y. So you 'll have y=cube root of (x/7)-9. Then solve for x. If I'm understanding your syntax (like jojo wrote), you'll have (y^3+9)/7=x. Swap x and y, and that's your answer (y=...)

OpenStudy (anonymous):

cuz if that is then I got \[(\sqrt[3]{49(x-63)})/7\]

OpenStudy (anonymous):

Oops! Mistake... hold on...

OpenStudy (anonymous):

which would make 49 7 and 63 9

OpenStudy (anonymous):

\[\sqrt[3]{7(x-9)}\]

OpenStudy (anonymous):

7(y^3+9)=x, not (y^3+9)/7=x. :) Jojo - don't try to take cube roots of those. Simply cube y and you get rid of the radical sign. Then add 9 to other side, multiply by 7.

OpenStudy (anonymous):

oops f−1(x)=7x3+63

OpenStudy (anonymous):

thats the answer sorry there didnt read the inverse

OpenStudy (anonymous):

Yeah, you were getting a massive answer! Couldn't figure out why! :)

OpenStudy (anonymous):

f^-1(x)= 7x^3+63

OpenStudy (anonymous):

sorry dkco

OpenStudy (anonymous):

No worries! I still think this is the answer: 7(x^3+9)=y. Remember this with inverses: if you want to check them, plug one equation into the other. They should come out as 1=1 or x=x. Good luck!

OpenStudy (anonymous):

I don't think I understand what you guys are doing. f^-1(x) = 21(x + 9) f^-1(x) = [7(x + 9)]^3 f^-1(x) = 7(x3 + 9) f^-1(x) = 7(x + 9)^3 these are my answers which are similar to what you're saying but not the same.

OpenStudy (isaiah.feynman):